# A Mixed-Order Phase Transition in Continuous-Variable Quantum Networks

### v1

@@ Abstract
The discontinuity around $\chi_{\text{th}}$ prompts us to investigate its critical nature, showing that the phase transition is, indeed, mixed order (SM, Sec. E). Below the critical threshold ($\chi<\chi_{\text{th}}$), we find that $\mathrm{X}_{\text{SC}}$ initially remains constant as a function of $l$, forming a plateau near $\mathrm{X}_{\text{SC}}\approx\mathrm{X}_{\text{SC}}^{+}$ [Fig. 1]. This plateau persists until $l$ reaches a finite value $l^{*}$, after which $\mathrm{X}_{\text{SC}}$ abruptly drops to zero. The same plateau is also observed in interdependent physical networks, signaling a dynamic characteristic of mixed-order phase transitions [46, *AD2022].
@@
By analogy with classical percolation [67], this value $l^{*}$ represents the characteristic length, signifying the typical distance within which two nodes remain connected. The log-log fitting [Fig. 1] further indicates that $l^{*}\sim\left|\chi-\chi_{\text{th}}\right|^{-1/2}$, suggesting that $l^{*}$ approaches infinity as $\chi\to\chi_{\text{th}}^{-}$, with a power law exponent $1/2$ that is equal to the chemical-space (“time”) thermal exponent, $z\nu$, as determined by finite-size scaling analysis [70]. Notably, $z\nu\approx 1/2$ differs from the classical and concurrence percolation theories, where $z\nu\approx 1$ for both [28, *OXSGBJ2022], indicating that the underlying dynamics of the NegPT is fundamentally unique and belongs to a different universality class.
@@
To identify the other critical exponent, $\beta$, note that akin to concurrence percolation, the series rule of the NegPT also has a “gauge redundancy” under the change of variable $\chi\to\chi^{x}$ [28, *OXSGBJ2022]. However, this redundancy no longer leads to arbitrariness in $\beta$. Indeed, for arbitrary $x$, we always have $\left(\mathrm{X}_{\text{SC}}\right)^{x}-\left(\mathrm{X}_{\text{SC}}^{+}\right)^{x}\simeq x\left(\mathrm{X}_{\text{SC}}^{+}\right)^{x-1}\left(\mathrm{X}_{\text{SC}}-\mathrm{X}_{\text{SC}}^{+}\right)\sim\left(\chi-\chi_{\text{th}}\right)^{1/2}$ (SM, Sec. E), indicating a fixed critical exponent $\beta=1/2$ [Fig. 1] independent of the gauge redundancy $x$.
@@
Mixed-order phase transitions also occur in classical interdependent percolation [71, *interdependent_network_2015], prompting a comparison to NegPT. Consider classical percolation in $M$ fully interdependent multilayer systems [73, *interdependent_mp19] of the Bethe lattice topology, and let $P_{\text{SC}}^{(1)}$ represent the sponge-crossing probability across a single branch [shaded orange in Fig. 1]. Per the excess-degree generating function interpretation [75], full interdependency requires that all branches in the $M$ layers must be simultaneously percolating in order to enable the percolation of the next-level branch [71, *interdependent_network_2015]. This consideration leads to a self-consistent equation (SM, Sec. F) that also features a mixed-order jump from 00 to $P_{\text{SC}}^{(1),+}$ at the critical threshold $p=p_{\text{th}}$. As $k\to\infty$, one has $P_{\text{SC}}^{(1),+}\to 0$, which leads to the expansion (SM, Sec. G)
@@
near the critical threshold. Analogously, for the NegPT, we find that the jump of the sponge-crossing ratio negativity across a single branch, $\mathrm{X}_{\text{SC}}^{(1),+}$, satisfies (SM, Sec. G)
@@
for $k>2$. Comparing Eqs. (3) and (4), we see that as $k\to\infty$, the NegPT mirrors classical percolation in a $(\mathrm{X}_{\text{SC}}^{(1),+})^{-2}$-layer interdependent Bethe lattice under the change of variable $\chi^{2}\to p$. In general, however, the two models are driven by fundamentally different mechanisms. While classical interdependent percolation requires an additional degree of freedom (i.e., the number of layers $M$) to induce a mixed-order transition, the NegPT does not require any such extra freedom.
@@
Generalized NegPT.—We also consider extending the series/parallel rules by introducing additional factors to Table 2, modifying the series rule to $\tanh r=\eta_{s}\tanh r_{1}\tanh r_{2}\dots$ and the parallel rule to $\sinh r=\eta_{p}\sinh r_{1}\cosh r_{2}\cosh r_{3}\dots$. While these generalized rules may not apply over the entire range of $r$ for arbitrary $\eta_{s}$ and $\eta_{p}$, we expect them to dominate near the critical threshold. Importantly, for the Bethe lattice, we find that a second-order phase transition is, indeed, forbidden when $\eta_{p}<\sqrt{k-1}$; while a mixed-order phase transition is forbidden when $\eta_{p}\geq\sqrt{k-1}$ (SM, Sec. A).
@@
What could be the possible physical origins of $\eta_{s}$ and $\eta_{p}$? First, the entanglement swapping may take a more general form [76, *GJ2002, *gaussian-povm_fm07], leading to $\eta_{s}\leq 1$ which reflects the generalization of homodyne measurement (SM, Sec. B). Second, the direct-product construction of the concentration operation (SM, Sec. B) appears to rely solely on the exponential tail of TMSVS in the Fock space, $\sim\chi^{n}|nn\rangle$. Therefore, we may consider generic CV states that are non-Gaussian at low photon numbers while retaining the exponential tail at high photon numbers. This could yield $\eta_{p}\leq 1$, reflecting the arbitrariness in the low-photon coefficients (SM, Sec. C). Notably, comparing the ranges $\eta_{s},\eta_{p}\leq 1$ with $\eta_{p}<\sqrt{k-1}$ in the Bethe lattice example suggests that mixed-order transitions are not ruled out and may, therefore, persist with generic CV operations and states.
@@
Instability of CV-based QN.—The identified abrupt transition distinguishes CV-based QNs from DV-based QNs, revealing potential instability inherent in the QN design. To illustrate this, note that in any realistic scenario, quantum states and their entanglement inevitably evolve over time due to environmental influences [79, 80, 81, 82, 83, 84, 85, 86]. As entanglement degradation is often an exponentially fast process [81, 82], we assume the ratio negativity $\chi(t)$ of each link follows an exponential decay, $\chi(t)\sim\exp{\left(-t/\tau\right)}$. This implies a decay rate proportional to the current negativity, $d\chi(t)/dt\sim-\tau^{-1}\chi(t)$, where the timescale $\tau$ characterizes the system’s noise and decoherence. Although this decay is a local process (as it occurs per link), its collective effect can rapidly compromise global functionality, driving $\mathrm{X}_{\text{SC}}\to 0$ particularly as $\chi(t)$ breaches $\chi_{\text{th}}$.
@@
Here, the first term represents decay of the ratio negativity over time and $u(t)$ signifies the feedback action applied after a delay $T_{0}$. The feedback signal $u(t)$ is determined by the observed deviation $\epsilon(t)$ (SM, Sec. V). In a physical system, for instance, $u(t)$ might correspond to increasing the pump laser power for down-conversion sources, boosting the squeezing as well as entanglement to compensate for ongoing losses [87, *down_conversion2010, *down_conversion2017, 80].
Our aim is to understand how the sharp, discontinuous nature of $\mathrm{X}_{\text{SC}}$ vs. $\chi$ [Fig. 1] undermines the feedback efficacy. Simulating an FOPTD model with identical $\tau$, $T_{0}$, and controller $u(t)$ on both DV- and CV-based Bethe-lattice QNs reveals starkly different behaviors. Under the concurrence percolation [28, *OXSGBJ2022], the sponge-crossing concurrence $C_{\text{SC}}$ for DV-based states recovers smoothly under feedback [Fig 2]. In contrast, the sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}$ for CV-based states, defined by its mixed-order transition, exhibits significant instability near the critical threshold $\chi_{\text{th}}$. Here, even minor dips can trigger repetitive performance collapse ($\mathrm{X}_{\text{SC}}\to 0$), resulting in unstable “on/off” oscillations using the same $u(t)$ [Fig 2]. This instability persists across a wide range of $u(t)$’s settings, making CV system stabilization substantially more demanding than for DV counterparts (SM, Sec. V), calling for a more careful feedback design for CV-based QNs.
Figure 2: Feedback stabilization of QN (Bethe lattice, $k=3$) against entanglement decay [Eq. (5)]. Under the same feedback control, 2 rapid stabilization is observed for the DV-based QN, whereas 2 the CV-based QN displays long-term “on/off” instability due to the abrupt drop in Fig. 1.
Discussion.—Despite the prevalence of mixed-order phase transitions [45, 46, *AD2022, 48, 49, 50, 51], the underlying mechanisms remain not fully understood. The traditional perspective of spinodal points [90] (or tipping points [91]) seems not applicable here, since our negativity percolation model—based solely on series and parallel rules—lacks a manifestation of hysteresis or metastability (which is also commonly absent in interdependent percolation [71, *interdependent_network_2015]). Consequently, the emergence of singularity may not necessarily be mean-field as for spinodal points [90]. This heralds future theoretical investigations into how the interplay of series and parallel rules can induce mixed-order transitions. Such studies could unveil novel phase transition landscapes potentially observable in quantum optical experiments implementing our G-G DET scheme.
@@
Acknowledgement.—Y.Z. (Zhao), K.H., and J.H. were supported by the National Natural Science Foundation of China under Grants No. 12271394 and No. 12071336. This work was supported by the Israel Ministry of Innovation, Science & Technology (grant number 01017980), the Israel Science Foundation (grant number 2830/23), the Binational Israel-China Science Foundation (grant number 3132/19), and the European Union under the Horizon Europe grant OMINO (grant number 101086321).
Supplemental Materials
Section I Gaussian-to-Gaussian Deterministic Entanglement Transmission (G-G DET)
A G-G entanglement swapping
@@
$g_{1}=\frac{2\sinh 2r_{1}}{\cosh 2r_{1}+\cosh 2r_{2}},$ (I.1)
$g_{4}=\frac{2\sinh 2r_{2}}{\cosh 2r_{1}+\cosh 2r_{2}},$ (I.2)
@@
$\tanh r=\tanh r_{1}\tanh r_{2}.$ (I.3)
This protocol can be extended to a one-dimensional chain, consisting of nodes $S$, $T$, and $N-1$ intermediate nodes $R_{1},\dots,R_{N-1}$ (Fig. S1), leading to the following theorem:
@@
$\tanh{r}=\prod_{j=1}^{N}\tanh{r_{j}}.$ (I.4)
@@
We may also express Eq. (I.4) in terms of the ratio negativity. Let $\chi_{j}\in[0,1]$ be the ratio negativity of each input TMSVS $|\psi^{r_{j}}\rangle$, and let $\chi\in[0,1]$ be the ratio negativity of the final output TMSVS $|\psi^{r}\rangle$. We define the series function
$\text{seri}(\chi_{1},\chi_{2},...):=\chi=\prod^{N}_{j=1}\chi_{j},$ (I.5)
which corresponds to the series rule for the G-G DET (Table 2).
Figure S1: Continuous-variable entanglement swapping. Given $N$ TMSVSs $|\psi^{r_{1}}\rangle,|\psi^{r_{2}}\rangle,\dots,|\psi^{r_{N}}\rangle$, the final state $|\psi^{r}\rangle$ between $S$ and $T$ derived by entanglement swapping is a TMSVS, where the squeezing parameter $r$ satisfies Eq. (I.4).
B G-G entanglement concentration
@@
Majorization between infinite-dimensional vectors.—Let $\vec{x}$ and $\vec{y}$ be positive (each element is a non-negative real number), infinite-dimensional vectors. $\vec{x}$ is majorized by $\vec{y}$, denoted as $\vec{x}\prec\vec{y}$, iff
$\Sigma_{j=1}^{k}x_{j}^{\downarrow}\leqslant\Sigma_{j=1}^{k}y_{j}^{\downarrow}$ (I.6)
holds for all positive integers $k$ and the equality holds for $k=\infty$. Here, $x_{j}^{\downarrow}$ ($y_{j}^{\downarrow}$) is the $j$th element of vector $\vec{x}$ ($\vec{y}$) arranged in descending order. Majorization theory implies that a pure bipartite state $|\psi\rangle$ can be deterministically transformed into another state $|\psi^{\prime}\rangle$ using LOCC iff $\vec{\lambda}\prec\vec{\lambda}^{\prime}$, where $\vec{\lambda}$ and $\vec{\lambda^{\prime}}$ are the Schmidt-value vectors of $|\psi\rangle$ and $|\psi^{\prime}\rangle$, with $\vec{\lambda}=(\lambda_{1},\lambda_{2},\cdots)^{\mathrm{T}}$ and $\vec{\lambda^{\prime}}=(\lambda_{1}^{\prime},\lambda_{2}^{\prime},\cdots)^{\mathrm{T}}$ [92] where $\sum_{j}\lambda_{j}=\sum_{j}\lambda_{j}^{\prime}=1$. For a TMSVS $|\psi^{r}\rangle$, its Schmidt-value vector $\vec{\lambda}(r)$ has entries $\lambda_{n}=\left(1-\tanh^{2}r\right)\tanh^{2n}(r)$. The majorization relation $\vec{x}\prec\vec{y}$ holds iff there exists a column-stochastic matrix $\mathbf{D}$ (whose column sum $=1$ for each column, and row sum $\leq 1$ for each row), such that $\vec{x}=\mathbf{D}\vec{y}$. Therefore, to construct deterministic LOCC for two TMSVSs with Schmidt-value vectors $\vec{x}$ and $\vec{y}$, respectively, conceptually one only needs to find a column-stochastic matrix $\mathbf{D}$ that satisfies $\vec{x}=\mathbf{D}\vec{y}$.
@@
$\mathbf{D}^{\mathcal{L}}=\begin{bmatrix}1&1-\eta&(1-\eta)^{2}&(1-\eta)^{3}&\dots\\ 0&{1\choose 1}\eta&{2\choose 1}\eta(1-\eta)&{3\choose 1}\eta(1-\eta)^{2}&\dots\\ 0&0&{2\choose 2}\eta^{2}&{3\choose 2}\eta^{2}(1-\eta)&\dots\\ 0&0&0&{3\choose 3}\eta^{3}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}.$ (I.7)
@@
$\mathbf{D}^{\mathcal{A}}=\begin{bmatrix}a&0&0&0&\dots\\ ab&a^{2}&0&0&\dots\\ ab^{2}&{2\choose 1}a^{2}b&a^{3}&0&\dots\\ ab^{3}&{3\choose 1}a^{2}b^{2}&{3\choose 2}a^{3}b&a^{4}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix},$ (I.8)
where $G>0$, $a=1/G$ and $b=1-a$.
@@
The introduction of these two matrices is particularly useful for TMSVS, since their tensor product $\mathbf{D}^{\mathcal{L}}\otimes\mathbf{D}^{\mathcal{A}}$ can transform the Schmidt-value vector of the tensor product of two TMSVSs, $\vec{\lambda}(r_{1}^{\prime})\otimes\vec{\lambda}(r_{2}^{\prime})$, into another vector of the same form $\vec{\lambda}(r_{1})\otimes\vec{\lambda}(r_{2})$, where $\tanh r_{1}^{\prime}=\tanh r_{1}/\sqrt{\eta+(1-\eta)\tanh^{2}r_{1}}$ and $\tanh r_{2}^{\prime}=\sqrt{1-G(1-\tanh^{2}r_{2})}$. The majorization relation $\vec{\lambda}(r_{1})\otimes\vec{\lambda}(r_{2})\prec\vec{\lambda}(r_{1}^{\prime})\otimes\vec{\lambda}(r_{2}^{\prime})$ holds if $\mathbf{D}^{\mathcal{L}}\otimes\mathbf{D}^{\mathcal{A}}$ is column-stochastic, which is true iff $\eta G\geqslant 1$. This gives rise to the following Lemma:
@@
$\frac{\sinh(r_{1}+r_{2})\pm\sinh(r_{1}-r_{2})}{\sinh(r_{1}^{\prime}+r_{2}^{\prime})\pm\sinh(r_{1}^{\prime}-r_{2}^{\prime})}\geqslant 1,$ (I.9)

### v2

@@ Results
The discontinuity around $\chi_{\text{th}}$ prompts us to investigate its critical nature, showing that the phase transition is, indeed, mixed order (SI, Sec. I E). Below the critical threshold ($\chi<\chi_{\text{th}}$), we find that $\mathrm{X}_{\text{SC}}$ initially remains constant as a function of $l$, forming a plateau near $\mathrm{X}_{\text{SC}}\approx\mathrm{X}_{\text{SC}}^{+}$ [Fig. 2(c)]. This plateau persists until $l$ reaches a finite value $l^{*}$, after which $\mathrm{X}_{\text{SC}}$ abruptly drops to zero. The same plateau is also observed in interdependent physical networks, signaling a dynamic characteristic of mixed-order phase transitions [98].
@@
By analogy with classical percolation [8], this value $l^{*}$ represents the characteristic length, signifying the typical distance within which two nodes remain connected. The log-log fitting [Fig. 2(d)] further indicates that $l^{*}\sim\left|\chi-\chi_{\text{th}}\right|^{-1/2}$, suggesting that $l^{*}$ approaches infinity as $\chi\to\chi_{\text{th}}^{-}$, with a power law exponent $1/2$ that is equal to the chemical-space (“time”) thermal exponent, $z\nu$, as determined by finite-size scaling analysis [44]. Notably, $z\nu\approx 1/2$ differs from the classical and concurrence percolation theories, where $z\nu\approx 1$ for both [66, 63], indicating that the underlying dynamics of the NegPT is fundamentally unique and belongs to a different universality class.
@@
To identify the other critical exponent, $\beta$, note that akin to concurrence percolation, the series rule of the NegPT also has a “gauge redundancy” under the change of variable $\chi\to\chi^{x}$ [66, 63]. However, this redundancy no longer leads to arbitrariness in $\beta$. Indeed, for arbitrary $x$, we always have $\left(\mathrm{X}_{\text{SC}}\right)^{x}-\left(\mathrm{X}_{\text{SC}}^{+}\right)^{x}\simeq x\left(\mathrm{X}_{\text{SC}}^{+}\right)^{x-1}\left(\mathrm{X}_{\text{SC}}-\mathrm{X}_{\text{SC}}^{+}\right)\sim\left(\chi-\chi_{\text{th}}\right)^{1/2}$ (SI, Sec. I E), indicating a fixed critical exponent $\beta=1/2$ [Fig. 2(b)] independent of the gauge redundancy $x$.
@@
Mixed-order phase transitions also occur in classical interdependent percolation [42, 48], prompting a comparison to NegPT. Consider classical percolation in $M$ fully interdependent multilayer systems [56, 69] of the Bethe lattice topology, and let $P_{\text{SC}}^{(1)}$ represent the sponge-crossing probability across a single branch [shaded orange in Fig. 2(a)]. Per the excess-degree generating function interpretation [86], full interdependency requires that all branches in the $M$ layers must be simultaneously percolating in order to enable the percolation of the next-level branch [42, 48]. This consideration leads to a self-consistent equation (SI, Sec. I F) that also features a mixed-order jump from 0 to $P_{\text{SC}}^{(1),+}$ at the critical threshold $p=p_{\text{th}}$. As $k\to\infty$, one has $P_{\text{SC}}^{(1),+}\to 0$, which leads to the expansion (SI, Sec. I G)
@@
near the critical threshold. Analogously, for the NegPT, we find that the jump of the sponge-crossing ratio negativity across a single branch, $\mathrm{X}_{\text{SC}}^{(1),+}$, satisfies (SI, Sec. I G)
@@
for $k>2$. Comparing Eqs. (3) and (4), we see that as $k\to\infty$, the NegPT mirrors classical percolation in an $(\mathrm{X}_{\text{SC}}^{(1),+})^{-2}$-layer interdependent Bethe lattice under the change of variable $\chi^{2}\to p$. In general, however, the two models are driven by fundamentally different mechanisms. While classical interdependent percolation requires an additional degree of freedom (i.e., the number of layers $M$) to induce a mixed-order transition, the NegPT does not require any such extra freedom.
To verify that the discontinuous nature of NegPT’s phase transition extends beyond the Bethe lattice, we also performed numerical simulations of entanglement percolation on a square lattice using the star-mesh transform [66]. As shown in Fig. 3(a), the curve drops abruptly already for finite side length $L$, suggesting a (weakly) discontinuous transition [40, 67] may indeed exist near the critical threshold $\chi_{\text{th}}\approx 0.715$. Moreover, we define the correlation length $\xi$ by $\chi\sim e^{-L/\xi}$ for $\chi<\chi_{\text{th}}$, measuring how fast $\chi$ approaches zero with respect to increasing length $L$ in the subcritical regime. Checking the scaling $\xi\sim|\chi-\chi_{\text{th}}|^{-\nu}$ yields an exponent $\nu=0.02(2)$ [Fig. 3(b)], a value that is distinct from $\nu\approx 1.3(3)$ for ConPT [66] and is close to zero, which is a clear signature of finite correlation length for discontinuous phase transitions [37].
.5 Generalized NegPT
@@
To examine how the mixed-order phase transition depends on the form of the series/parallel rules in Fig. 1(c), we consider generalizing these rules. Specifically, we introduce two constant factors, $\eta_{s}$ and $\eta_{p}$, modifying the series rule to $\tanh r=\eta_{s}\tanh r_{1}\tanh r_{2}\dots$ and the parallel rule to $\sinh r=\eta_{p}\sinh r_{1}\cosh r_{2}\cosh r_{3}\dots$. These generalized rules may not be physically valid across the entire parameter range for arbitrary $\eta_{s}$ and $\eta_{p}$. However, our quantitative analysis is restricted to the region near the critical threshold, where these forms serve as a valid local approximation. On the Bethe lattice, we find that the type of the phase transition depends on $\eta_{p}$. Specifically, when $\eta_{p}<\sqrt{k-1}$, only a mixed-order phase transition can exist; when $\eta_{p}\geq\sqrt{k-1}$, only a continuous phase transition can exist (SI, Sec. II A).
@@
What could be the possible physical origins of $\eta_{s}$ and $\eta_{p}$? First, the entanglement swapping may take a more general form [39, 43, 38], leading to $\eta_{s}\leq 1$ which reflects the generalization of homodyne measurement (SI, Sec. II B). Second, the direct-product construction of the concentration operation (see Methods) appears to rely solely on the exponential tail of TMSVS in the Fock space, $\sim\chi^{n}|nn\rangle$. Therefore, we may consider generic CV states that are non-Gaussian at low photon numbers while retaining the exponential tail at high photon numbers. This could yield $\eta_{p}\leq 1$, reflecting the arbitrariness in the low-photon coefficients (SI, Sec. II C). Interestingly, comparing this with the $\eta_{p}<\sqrt{k-1}$ criterion in the Bethe lattice example suggests that only mixed-order transitions can be observed with generic CV operations as in our settings.
Figure 3: Entanglement percolation in two-dimensional square lattices. (a) $\text{X}_{\text{SC}}$ for square lattices with different side length $L$. (b) Scaling of the correlation length $\xi$ near the critical threshold $\chi_{\text{th}}\approx 0.715$ follows $\xi\sim|\chi-\chi_{\text{th}}|^{-\nu}$, with a fitted critical exponent $\nu=0.02\pm 0.02$.
.6 Instability under feedback control
@@
The identified abrupt transition distinguishes CV-based QNs from DV-based QNs, revealing potential instability inherent in the QN design. To illustrate this, note that in any realistic scenario, quantum states and their entanglement inevitably evolve over time due to environmental influences [87, 2, 94, 27, 89, 46, 78, 5]. As entanglement degradation is often an exponentially fast process [94, 27], we assume the ratio negativity $\chi(t)$ of each link follows an exponential decay, $\chi(t)\sim\exp{\left(-t/\tau\right)}$. This implies a decay rate proportional to the current negativity, $d\chi(t)/dt\sim-\tau^{-1}\chi(t)$, where the timescale $\tau$ characterizes the system’s noise and decoherence. Although this decay is a local process (as it occurs per link), its collective effect can rapidly compromise global functionality, driving $\mathrm{X}_{\text{SC}}\to 0$ particularly as $\chi(t)$ breaches $\chi_{\text{th}}$.
@@
Here, the first term represents decay of the ratio negativity over time and $u(t)$ signifies the feedback action applied after a delay $T_{0}$. The feedback signal $u(t)$ is determined by the observed deviation $\epsilon(t)$ (SI, Sec. III A). In a physical system, for instance, $u(t)$ might correspond to increasing the pump laser power for down-conversion sources, boosting the squeezing as well as entanglement to compensate for ongoing losses [58, 7, 28, 2].
Figure 4: Feedback stabilization of QN against entanglement decay. (a) Under the same feedback control [Eq. (5)], the DV-based QN ($k=3$ Bethe lattice) exhibits rapid stabilization; (b) whereas the CV-based QN exhibits long-term “on/off” instability, a direct result of the abrupt drop in Fig. 2(b).
Our aim is to understand how the sharp, discontinuous nature of $\mathrm{X}_{\text{SC}}$ vs. $\chi$ [Fig. 2(b)] undermines the feedback efficacy. Simulating an FOPTD model with identical $\tau$, $T_{0}$, and controller $u(t)$ on both DV- and CV-based Bethe-lattice QNs reveals starkly different behaviors. Under the concurrence percolation [66, 63], the sponge-crossing concurrence $C_{\text{SC}}(t)$ for DV-based states recovers smoothly under feedback [Fig 4(a)]. In contrast, the sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}(t)$ for CV-based states, dictated by its mixed-order transition, exhibits unstable “on/off” oscillations using the same $u(t)$ [Fig. 4(b)].
Crucially, this instability is not a result of the limitation of the FOPTD model’s capability. Indeed, under Eq. (5), $\chi(t)$ is successfully stabilized within a narrow range and only undergoes small perturbations near the critical threshold $\chi_{\text{th}}\approx 0.866$ (SI, Sec. III B). Therefore, the instability of $\mathrm{X}_{\text{SC}}$ comes directly from the sharp nature of $\mathrm{X}_{\text{SC}}$ vs. $\chi$ [Fig. 2(b)]. We also find that the instability persists across a wide range settings for $u(t)$ (SI, Sec. III C), indicating that the CV system stabilization is substantially more demanding than for DV counterparts, calling for a more careful feedback design for CV-based QNs.
Discussion
Despite the rich studies on mixed-order phase transitions [71, 98, 14, 77, 92, 3], the underlying mechanisms remain not fully understood. The traditional perspective of spinodal points [57] seems not applicable here, since our negativity percolation model—based solely on series and parallel rules—lacks a manifestation of hysteresis or metastability (which is also commonly absent in interdependent percolation [42, 48]). Consequently, the emergence of singularity may not necessarily be described by a partition function as for spinodal points [57]. This heralds future theoretical investigations into how the interplay of series and parallel rules can induce mixed-order transitions. Such studies could unveil novel phase transition landscapes potentially observable in quantum optical experiments implementing our G-G DET scheme.
@@ Discussion
Methods
This section presents the foundational tools for studying negativity percolation on Gaussian quantum networks. Further details are provided in the Supplementary Information.
.1 G-G entanglement swapping
@@ Methods
$g_{1}=\frac{2\sinh 2r_{1}}{\cosh 2r_{1}+\cosh 2r_{2}},$ (6)
$g_{4}=\frac{2\sinh 2r_{2}}{\cosh 2r_{1}+\cosh 2r_{2}},$ (7)
@@
$\tanh r=\tanh r_{1}\tanh r_{2}.$ (8)
This protocol can be extended to a one-dimensional chain, consisting of nodes $S$, $T$, and N1N-1 intermediate nodes $R_{1},\dots,R_{N-1}$ (Fig. 5), leading to the following theorem:
@@
$\tanh{r}=\prod_{j=1}^{N}\tanh{r_{j}}.$ (9)
@@
We may also express Eq. (9) in terms of the ratio negativity. Let $\chi_{j}\in[0,1]$ be the ratio negativity of each input TMSVS $|\psi^{r_{j}}\rangle$, and let $\chi\in[0,1]$ be the ratio negativity of the final output TMSVS $|\psi^{r}\rangle$. We define the series function
$\text{seri}(\chi_{1},\chi_{2},...):=\chi=\prod^{N}_{j=1}\chi_{j},$ (10)
which corresponds to the series rule for the G-G DET.
Figure 5: Continuous-variable entanglement swapping. Given $N$ TMSVSs $|\psi^{r_{1}}\rangle,|\psi^{r_{2}}\rangle,\dots,|\psi^{r_{N}}\rangle$, the final state $|\psi^{r}\rangle$ between $S$ and $T$ derived by entanglement swapping is a TMSVS, where the squeezing parameter $r$ satisfies Eq. (9).
.2 G-G entanglement concentration
@@
We now introduce the majorization theory between infinite-dimensional vectors. Let $\vec{x}$ and $\vec{y}$ be positive (each element is a non-negative real number), infinite-dimensional vectors. $\vec{x}$ is majorized by $\vec{y}$, denoted as $\vec{x}\prec\vec{y}$, iff
$\Sigma_{j=1}^{k}x_{j}^{\downarrow}\leq\Sigma_{j=1}^{k}y_{j}^{\downarrow}$ (11)
holds for all positive integers $k$ and the equality holds for $k=\infty$. Here, $x_{j}^{\downarrow}$ ($y_{j}^{\downarrow}$) is the $j$th element of vector $\vec{x}$ ($\vec{y}$) arranged in descending order. Majorization theory implies that a bipartite pure state $|\psi\rangle$ can be deterministically transformed into another state $|\psi^{\prime}\rangle$ using LOCC iff $\vec{\lambda}\prec\vec{\lambda}^{\prime}$, where $\vec{\lambda}$ and $\vec{\lambda^{\prime}}$ are the Schmidt-value vectors of $|\psi\rangle$ and $|\psi^{\prime}\rangle$, with $\vec{\lambda}=(\lambda_{1},\lambda_{2},\cdots)^{\mathrm{T}}$ and $\vec{\lambda^{\prime}}=(\lambda_{1}^{\prime},\lambda_{2}^{\prime},\cdots)^{\mathrm{T}}$ [73] where $\sum_{j}\lambda_{j}=\sum_{j}\lambda_{j}^{\prime}=1$. For a TMSVS $|\psi^{r}\rangle$, its Schmidt-value vector $\vec{\lambda}(r)$ has entries $\lambda_{n}=\left(1-\tanh^{2}r\right)\tanh^{2n}(r)$. The majorization relation $\vec{x}\prec\vec{y}$ holds iff there exists a column-stochastic matrix $\mathbf{D}$ (whose column sum =1=1 for each column, and row sum $\leq 1$ for each row), such that $\vec{x}=\mathbf{D}\vec{y}$. Therefore, to construct deterministic LOCC for two TMSVSs with Schmidt-value vectors $\vec{x}$ and $\vec{y}$, respectively, conceptually one only needs to find a column-stochastic matrix $\mathbf{D}$ that satisfies $\vec{x}=\mathbf{D}\vec{y}$.
@@
$\mathbf{D}^{\mathcal{L}}=\begin{bmatrix}1&1-\eta&(1-\eta)^{2}&(1-\eta)^{3}&\dots\\ 0&{1\choose 1}\eta&{2\choose 1}\eta(1-\eta)&{3\choose 1}\eta(1-\eta)^{2}&\dots\\ 0&0&{2\choose 2}\eta^{2}&{3\choose 2}\eta^{2}(1-\eta)&\dots\\ 0&0&0&{3\choose 3}\eta^{3}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}$. (12)
@@
$\mathbf{D}^{\mathcal{A}}=\begin{bmatrix}a&0&0&0&\dots\\ ab&a^{2}&0&0&\dots\\ ab^{2}&{2\choose 1}a^{2}b&a^{3}&0&\dots\\ ab^{3}&{3\choose 1}a^{2}b^{2}&{3\choose 2}a^{3}b&a^{4}&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}$, (13)
where $G>0$, $a=1/G$ and b=1ab=1-a.
@@
The introduction of these two matrices is particularly useful for TMSVS, since their tensor product $\mathbf{D}^{\mathcal{L}}\otimes\mathbf{D}^{\mathcal{A}}$ can transform the Schmidt-value vector of the tensor product of two TMSVSs, $\vec{\lambda}(r_{1}^{\prime})\otimes\vec{\lambda}(r_{2}^{\prime})$, into another vector of the same form $\vec{\lambda}(r_{1})\otimes\vec{\lambda}(r_{2})$, where $\tanh r_{1}^{\prime}=\tanh r_{1}/\sqrt{\eta+(1-\eta)\tanh^{2}r_{1}}$ and $\tanh r_{2}^{\prime}=\sqrt{1-G(1-\tanh^{2}r_{2})}$. The majorization relation $\vec{\lambda}(r_{1})\otimes\vec{\lambda}(r_{2})\prec\vec{\lambda}(r_{1}^{\prime})\otimes\vec{\lambda}(r_{2}^{\prime})$ holds if $\mathbf{D}^{\mathcal{L}}\otimes\mathbf{D}^{\mathcal{A}}$ is column-stochastic, which is true iff $\eta G\geq 1$. This gives rise to the following Lemma:
@@
$\frac{\sinh(r_{1}+r_{2})\pm\sinh(r_{1}-r_{2})}{\sinh(r_{1}^{\prime}+r_{2}^{\prime})\pm\sinh(r_{1}^{\prime}-r_{2}^{\prime})}\geq 1,$ (14)